OpenStudy (anonymous):

show that cos(x)> 1-(x^2/2)

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (tkhunny):

How about proving that it is false?! Try x = 0.

OpenStudy (tkhunny):

Do we get the Taylor Series?

OpenStudy (anonymous):

Since this is differentiable for all R, let's apply the Mean Value Theorem to f for all x between 0 and t: f(t) - f(0) = f '(c) (t - 0) for some c between 0 and t ==> cos t - 1 + t^2/2 = (c - sin c) * t. However, c ≥ sin c for all c ≥ 0. ==> (c - sin c) * t ≥ 0 for t ≥ 0. Similarly, c ≤ sin c for all c < 0 (by using the above inequality and multiplying both sides by -1) ==> (c - sin c) * t ≥ 0 for t < 0. Therefore, cos t - 1 + t^2/2 ≥ 0 for all t ==> cos t ≥ 1 - t^2/2 for all t. I hope this helps!

OpenStudy (tkhunny):

First, we must establish that \(\ge\) was intended and not \(>\).